OpenStudy (anonymous):

What is the slope of the line? (4,9) and (2,9)

OpenStudy (anonymous):

I really need the help please

OpenStudy (anonymous):

Divide the vertical distance by the horizontal distance. Vertical distance: 9 - 9 = 0. Horizontal distance: 2 - (4) = -2 0 / (-2) = 0 The line's slope is zero.

OpenStudy (anonymous):

Oh my gosh thank you so so much

OpenStudy (anonymous):

Welcome!